OpenStudy (anonymous):

Hello, I need help with telling whether the side lengths form a pythagorean triple, and to find the unknown side lengths, Please Help

Directrix (directrix):

Do you know any of the "basic" PYthagorean triples?

Directrix (directrix):

Here is a chart of some of them to give an idea of what they are.

OpenStudy (freckles):

\[(a,b, \sqrt{a^2+b^2}) \text{ is a pythagorean triple if } a, b , \text{ and } \sqrt{a^2+b^2} \text{ are integers } \\ \text{ where } \sqrt{a^2+b^2}>a \text{ and } b\]